# Goal-Oriented Multiple Access Connectivity for Networked Intelligent Systems

### v1

@@ III-B Effective Resource Consumption (ERC)
The ERC feature allows us to model the resource consumption for communicating the content as a function of its usefulness. Having time-slotted channels with shared frequency band, power is the main candidate resource to adapt. Therefore, we define the transmission power of the $k$-th ISA at the $j$-th time slot as $\rho_{k,j}^{(n)}=f_{\rho}(v_{k,i}^{(n)})$, in the form of a non-decreasing function $f_{\rho}:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$ of the meta-value assigned to $x_{i}^{(n)}$ at the $j$-th slot. In this regard, we can write
${\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{2}\prod_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}\mathcal{S}_{n}$ (8)
where $\text{where~{}}\mathcal{F}_{2}=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\sum_{k\in\mathcal{K}_{n}}h\!\left(\alpha_{k,n}\dfrac{f_{\rho}(v_{k,j}^{(n)})}{v_{k,j}^{(n)}}\right)\right]_{n\leftarrow\mathcal{A}_{j}}\!,$
$h:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$, and $\mathcal{S}_{n}=1-\mathcal{E}_{n}$ is the probability of correctly delivering the $n$-th attribute to the required number of NMAs.
III-C Effective Utility of Content (EUoC)
The EUoC feature shapes the utility (usefulness) of the contents in satisfying the goal once the actual event they are representing is correctly reconstructed at the NMAs. Thereby, the EUoC at each service period depends on the sum value of the communicated attributes at that period, modeled as follows
${\rm{EUoC}}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\prod_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}\mathcal{S}_{n}.$ (9)
@@ IV Multiple Access Design
We aim to design a multiple access scheme via obtaining the set of optimal activation probabilities $\boldsymbol{\alpha}^{*}=[\alpha_{k,n}^{*}]_{K\times N}$, maximizing the GoE $\mathcal{G}$ in the system based on the goal. Next, we find the threshold criteria for decision-making and propose a self-decision scheme under all content acquisition schemes.
@@ IV-A Problem Formulation
With regard to the definition of $\mathcal{G}$ in Section III, we formulate the following optimization problem
$\mathcal{P}:~{}$ $\underset{\boldsymbol{\alpha},\mathbf{W}}{\text{max}}~{}\mathcal{G}=\varphi\Big{(}\mathbf{W};g_{1}\big{(}I_{1}\big{)},g_{2}\big{(}I_{2}\big{)},...,g_{c}\big{(}I_{c}\big{)}\Big{)}$
@@
where $0\prec\mathbf{W}=[{w}_{r}]_{c\times 1}\preceq 1$ is a weight vector. We reform the GoE optimization in $\mathcal{P}$ to a problem of minimizing the weighted sum of the EDE and ERC features, subject to guaranteeing that the EUoC exceeds the minimum desirable effectiveness $U_{\rm GoE}$ at each service period. As transmissions are performed independently among the service periods, it is sufficient to analyze the effectiveness problem for an arbitrary $t$-th period. Then we can apply the same access design for other periods, depending on $\widehat{\mathcal{A}}_{t}$, $\forall t$. Finally, the problem becomes
$\widehat{\mathcal{P}}:~{}$ $\underset{\boldsymbol{\alpha},\mathbf{W}}{\text{min}}\left[w_{1}g_{1}\big{(}{\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})\big{)}\!+\!w_{2}g_{2}\big{(}{\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})\big{)}\right]$
@@
[…]
where $g_{r}:\mathbb{R}_{0}^{+}\to\mathbb{R}$, for $r\in\{1,2,3\}$, indicates a positive non-decreasing differentiable function.
@@ IV-B Problem Solution
Inserting (2), (8), and (9) into (IV-A), we define the Lagrange function $\mathcal{L}$ for $\widehat{\mathcal{P}}$, and we can write the Karush-Kuhn-Tucker (KKT) conditions, as follows
$\dfrac{\partial\mathcal{L}}{\partial\alpha_{k,n}}$ $=w_{1}\Bigg{[}\mathcal{F}_{1}\!\left(\dfrac{\partial P_{{\rm e},n}}{\partial\alpha_{k,n}}\right)\!+\!P_{{\rm e},n}\!\left(\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\right)\!\Bigg{]}g_{1}^{\prime}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)$
@@
where $\overline{\mathcal{S}}_{n}=\!\!\prod_{n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus{\{n\}}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid-1}}}\!\mathcal{S}_{n^{\prime}}$. Moreover, we have
$\dfrac{\partial\mathcal{L}}{\partial w_{1}}$ $=g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)\!+\mu_{0}+\mu_{1}=0,$ (13)
@@
In (IV-B)–(14), $\eta\geq 0$, $\mu_{r}\geq 0$, $\forall r$, and $\lambda_{k,n}\geq 0$, $\forall n\in\widehat{\mathcal{A}}_{t}$, $\forall k\in\mathcal{K}_{n}$, denote Lagrange multipliers.
@@
$\mathcal{H}_{1}:\eta\Big{(}g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!-U_{\rm GoE}\Big{)}=0,$
@@
Each of $\mathcal{H}_{1}$ to $\mathcal{H}_{3}$ holds under two conditions: In $\mathcal{H}_{3}$, we consider (i) $\lambda_{k,n}=0$, hence $\alpha_{k,n}<1$; or (ii) $\alpha_{k,n}=1$, thus $\lambda_{k,n}>0$. Under condition (ii), the ISAs are active at all time slots, which is not optimal according to (4) and (5). Thus, we reach $\lambda_{k,n}=0$, $\forall k,n$. In $\mathcal{H}_{2}$, we assume (i) $\mu_{r}=0$, then $w_{r}<1$; or (ii) $w_{r}=1$, hence $\mu_{r}>0$. Condition (ii) leads one of the weighting variables to become zero, which is incorrect referring to $\widehat{\mathcal{P}}$. Thus, we have $\mu_{r}=0$, $\forall r$. Given this, we arrive
$\mathcal{F}_{2}=\dfrac{1}{\mathcal{S}_{n}\overline{\mathcal{S}}_{n}}g_{2}^{-1}\!\!\left[g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n\in\widehat{\mathcal{A}}_{t}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid}}}P_{{\rm e},n}\right)\!\right]\!.$ (15)
Also, for $\mathcal{H}_{1}$, we have (i) $\eta=0$, and $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)>U_{\rm GoE}$; or (ii) $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)=U_{\rm GoE}$, and $\eta>0$. Applying (15) into (IV-B), condition (i) results in negative probabilities. Thus, condition (ii) is true with $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)=U_{\rm GoE}$. From (IV-B), we obtain
$P_{{\rm e},n}=\left[\dfrac{g_{1}^{-1}\!\!\left(\dfrac{\eta\mathcal{S}_{n}\overline{\mathcal{S}}_{n}U_{\rm GoE}}{2w_{1}}\right)}{\mathcal{F}_{1}\sum_{n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus{\{n\}}}^{{{\mid\widehat{\mathcal{A}}_{t}\mid-1}}}P_{{\rm e},n^{\prime}}}\right]^{\!1/2}\!.$ (16)
@@
Input: Known parameters $M$, $L$, $p$, $\widehat{\mathcal{A}}_{t}$, $\mathcal{K}_{n}$, $v_{k,j}^{(n)}$, $\forall j,k$, $\chi_{k,n}$, $\gamma^{(\mathcal{C},\mathcal{I})}_{k,n}$, $\forall n,k$, $U_{\rm GoE}$, $\boldsymbol{\varepsilon_{1}}$, and $\boldsymbol{\varepsilon_{2}}$. Initial parameters $\eta^{(0)}\to 0$, $\boldsymbol{\alpha}^{(0)}=\mathbf{0}_{K\times|\widehat{\mathcal{A}}_{t}|}$, and $\mathbf{W}=[0.5,0.5]^{T}$. Forms of $g_{r}$, $\forall r$, $h$, and $f_{\rho}$.
@@ 1 Iteration $j$: $\triangleright$ Outer loop
2 if conditions $\mathcal{H}_{4}$ to $\mathcal{H}_{6}$ are met. then goto 3. Iteration $i$: $\triangleright$ Inner loop
@@
5 Derive $\alpha_{k,n}^{*}$ from (4), (5), (1), (16), and given $\alpha_{k^{\prime},n}$, $\forall n,k^{\prime}\in\mathcal{K}_{n}\setminus\{k\}$; update and save $\boldsymbol{\alpha}^{(i)}$.
@@
7 if $\Big{|}\boldsymbol{\alpha}^{(i)}\!-\!\boldsymbol{\alpha}^{(i-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{1}}$ then set $i\!=\!i\!+\!1$, and goto 3. Compute $w_{1}^{(j)}$ and $w_{2}^{(j)}$; save $\mathbf{W}^{(j)}$.
if $\Big{|}\mathbf{W}^{(j)}\!-\!\mathbf{W}^{(j-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{2}}$, $\mathbf{W}^{(j)}\preceq 0$, $\mathbf{W}^{(j)}\succeq 1$, or $\exists_{k,n}\alpha_{k,n}^{(i)}<0$ then step up $\eta^{(j)}$, $j\!=\!j\!+\!1$, and goto 1. Return $\boldsymbol{\alpha}^{*}=\boldsymbol{\alpha}^{(i)}$ and $\mathbf{W}^{*}=\mathbf{W}^{(j)}$.
Algorithm 1 Solution for deriving $\boldsymbol{\alpha}^{*}$ and $\mathbf{W}^{*}$ at the $t$-th service period.
@@ IV-C Analysis for Large $I_{n}$
We consider the attribute $\mathcal{X}_{n}$, $\forall n$, has large enough state space, i.e., $I_{n}\gg 1$. In this regime, we achieve $\underset{L\to\infty}{\lim}P_{{\rm e},n}=\frac{1}{2}$ from (4). Thus, using (16), $\alpha^{*}_{k,n}$, $\forall k,n$, is derived as follows
$\alpha_{k,n}^{*}=\dfrac{\eta\overline{\mathcal{S}}_{n}U_{\rm GoE}\Big{(}1\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}-2w_{1}g_{1}\!\!\left(\dfrac{\mathcal{F}_{1}}{2^{(|\widehat{\mathcal{A}}_{t}|+1)}}\right)}{\eta\overline{\mathcal{S}}_{n}U_{\rm GoE}\Big{(}{\mathcal{E}}_{n}^{(k)}\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}\!}$ (18)
where we rewrite (1) and define ${\mathcal{E}}_{n}^{(k)}$ and ^$\widehat{\mathcal{E}}_{n}^{(k)}$ as below
${\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}_{\ell}}\tilde{\mathcal{E}}_{n,\ell}}{\alpha_{k,n}},~{}\,\widehat{\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}^{\prime}_{\ell}}\tilde{\mathcal{E}}_{n,\ell}}{1-\alpha_{k,n}}.$
From (18), we infer that $\alpha_{k,n}^{*}$ increases by (i) decreasing the average usefulness of contents to tackle the high contamination of useful contents over the channels, (ii) enlarging the set of queried attributes, and (iii) increasing the $U_{\rm GoE}$ constraint.

### v2

@@ III-B Effective Resource Consumption (ERC)
The ERC feature allows us to model the resource consumption for communicating updates as a function of its usefulness. Having time-slotted channels with shared frequency band, power is the main candidate resource to adapt. Therefore, we define the transmission power of the $k$-th ISA at the $j$-th time slot as $\rho_{k,j}^{(n)}=f_{\rho}(v_{k,i}^{(n)})$, in the form of a non-decreasing function $f_{\rho}:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$ of the meta value assigned to $x_{i}^{(n)}$ at the $j$-th slot. In this regard, we can write
${\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{2}\prod_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}\mathcal{S}_{n}$ (7)
where $\text{where~{}}\mathcal{F}_{2}=\sum_{j=1}^{{\mid\widehat{\mathcal{A}}_{t}\mid}}\left[\sum_{k\in\mathcal{K}_{n}}h\!\left(\alpha_{k,n}\dfrac{f_{\rho}(v_{k,j}^{(n)})}{v_{k,j}^{(n)}}\right)\right]_{n:\mathcal{X}_{n}=\mathcal{X}_{j}}\!,$
$h:\mathbb{R}^{+}_{0}\to\mathbb{R}^{+}_{0}$ is a non-decreasing differentiable function, and $\mathcal{S}_{n}=1-\mathcal{E}_{n}$ denotes the probability of correctly delivering the $n$-th attribute to the required number of NMAs.
III-C Effective Utility of Updates (EUU)
The EUU feature shapes the utility (usefulness) of updates in satisfying the goal once the actual event they are representing is correctly reconstructed at the NMAs. Thereby, the EUU at each service interval depends on the sum value of the communicated attributes at that period, modeled as follows
${\rm{EUU}}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})=\mathcal{F}_{1}\prod_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}\mathcal{S}_{n}.$ (8)
@@ IV Multiple Access Design
We aim to design a multiple access scheme via obtaining the set of optimal activation probabilities $\boldsymbol{\alpha}^{*}=[\alpha_{k,n}^{*}]_{K\times N}$, maximizing the GoE in the system based on the goal. Next, we find the threshold criteria for decision-making and propose a self-decision scheme under all update acquisition schemes.
@@ IV-A Problem Formulation
Based on Section III, we formulate the problem as follows
$\mathcal{P}_{1}:~{}\underset{\boldsymbol{\alpha},\mathbf{W}}{\text{max}}~{}{\rm GoE}=\varphi\left(\mathbf{W}^{T};g_{1}(f_{1}),\dots,g_{R}(f_{R})\right)$
@@
Without loss of generality, we transform the GoE optimization in $\mathcal{P}_{1}$ to a problem of minimizing the weighted sum of the EDE and ERC features, named $\mathcal{G}$, subject to guaranteeing that the EUU exceeds a minimum value denoted by ${\rm{EUU}}_{\rm min}$ at each service interval. As transmissions are performed independently among the service intervals, it is sufficient to analyze the effectiveness problem for an arbitrary $t$-th period. Then, we can apply the same access design for the other periods, depending on $\widehat{\mathcal{A}}_{t}$, $\forall t$. Therefore, we can write
$\mathcal{P}_{2}:~{}\underset{0\preceq\boldsymbol{\alpha}\preceq 1,\,0\prec\mathbf{W}\preceq 1}{\text{min}}~{}\mathcal{G}=\sum_{r=1}^{2}w_{r}g_{r}(f_{r})$
@@
where $f_{1}={\rm EDE}(E_{t},\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})$, and $f_{2}={\rm ERC}(\widehat{\mathcal{A}}_{t};\boldsymbol{\alpha},\mathbf{v})$.
@@ IV-B Problem Solution
Inserting (2), (7), and (8) into (IV-A), we define the Lagrange function $\mathcal{L}$ for $\mathcal{P}_{2}$, and we can write the Karush-Kuhn-Tucker (KKT) conditions, as follows
$\dfrac{\partial\mathcal{L}}{\partial\alpha_{k,n}}=$
$~{}~{}~{}w_{1}\Bigg{[}\mathcal{F}_{1}\!\left(\dfrac{\partial P_{{\rm e},n}}{\partial\alpha_{k,n}}\right)\!+\!P_{{\rm e},n}\!\left(\dfrac{\partial\mathcal{F}_{1}}{\partial\alpha_{k,n}}\right)\!\Bigg{]}g_{1}^{\prime}\!\!\left(\mathcal{F}_{1}\sum_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}P_{{\rm e},n}\right)$
@@
where $\overline{\mathcal{S}}_{n}=\!\!\prod_{n^{\prime}\in\widehat{\mathcal{A}}_{t}\setminus{\{\mathcal{X}_{n}\}}}\!\mathcal{S}_{n^{\prime}}$. Moreover, we have
$\dfrac{\partial\mathcal{L}}{\partial w_{1}}$ $=g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}P_{{\rm e},n}\right)\!+\mu_{0}+\mu_{1}=0,$ (12)
@@
In (IV-B)–(13), $\eta\geq 0$, $\mu_{r}\geq 0$, $\forall r$, and $\lambda_{k,n}\geq 0$, $\forall n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}$, $\forall k\in\mathcal{K}_{n}$, denote Lagrange multipliers.
@@
$\mathcal{H}_{1}:\eta\Big{(}g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)\!-{\rm{EUU}}_{\rm min}\Big{)}=0,$
@@
Each of $\mathcal{H}_{1}$ to $\mathcal{H}_{3}$ holds under two conditions: In $\mathcal{H}_{3}$, we consider (i) $\lambda_{k,n}=0$, hence $\alpha_{k,n}<1$; or (ii) $\alpha_{k,n}=1$, thus $\lambda_{k,n}>0$. Under condition (ii), the ISAs are active at all time slots, which is not optimal according to (4). Thus, we reach $\lambda_{k,n}=0$, $\forall k,n$. In $\mathcal{H}_{2}$, we assume (i) $\mu_{r}=0$, then $w_{r}<1$; or (ii) $w_{r}=1$, hence $\mu_{r}>0$. Condition (ii) leads one of the weighting variables to become zero, which is incorrect referring to $\mathcal{P}_{2}$. Thus, we have $\mu_{r}=0$, $\forall r$. Given this, we arrive
$\mathcal{F}_{2}=\dfrac{1}{\mathcal{S}_{n}\overline{\mathcal{S}}_{n}}g_{2}^{-1}\!\!\left[g_{1}\!\!\left(\mathcal{F}_{1}\sum_{n:\mathcal{X}_{n}\in\widehat{\mathcal{A}}_{t}}P_{{\rm e},n}\right)\!\right]\!.$ (14)
Also, for $\mathcal{H}_{1}$, we have (i) $\eta=0$, and $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)>{\rm{EUU}}_{\rm min}$; or (ii) $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)={\rm{EUU}}_{\rm min}$, and $\eta>0$. Applying (14) into (IV-B), condition (i) results in negative probabilities. Thus, condition (ii) is true with $g_{3}\!\left(\mathcal{F}_{1}\mathcal{S}_{n}\overline{\mathcal{S}}_{n}\right)={\rm{EUU}}_{\rm min}$. From (IV-B), we obtain
$P_{{\rm e},n}=\left[\dfrac{g_{1}^{-1}\!\!\left(\dfrac{\eta\mathcal{S}_{n}\overline{\mathcal{S}}_{n}{\rm{EUU}}_{\rm min}}{2w_{1}}\right)}{\mathcal{F}_{1}\sum_{n^{\prime}:\mathcal{X}_{n^{\prime}}\in\widehat{\mathcal{A}}_{t}\setminus{\{\mathcal{X}_{n}\}}}P_{{\rm e},n^{\prime}}}\right]^{\!1/2}\!.$ (15)
@@
Input: Known parameters $M$, $L$, $p$, $\widehat{\mathcal{A}}_{t}$, $\mathcal{K}_{n}$, $v_{k,j}^{(n)}$, $\forall j,k$, $q_{k,n}$, $\gamma_{k,m,n}^{(\mathcal{C},\mathcal{I})}$, $\forall n,k$, ${\rm{EUU}}_{\rm min}$, $\boldsymbol{\varepsilon_{1}}$, and $\boldsymbol{\varepsilon_{2}}$. Initial parameters $\eta^{(0)}\leftarrow 0$, $\boldsymbol{\alpha}^{(0)}=\mathbf{0}_{K\times|\widehat{\mathcal{A}}_{t}|}$, and $\mathbf{W}=[0.5,0.5]^{T}$. Forms of $g_{r}$, $\forall r$, $h$, and $f_{\rho}$.
@@ 1 Iteration $j$: $\triangleright$ Outer loop
2 if conditions $\mathcal{H}_{4}$ to $\mathcal{H}_{6}$ are met. then goto 3.
3 Iteration $i$: $\triangleright$ Inner loop
@@
6 Derive $\alpha_{k,n}^{*}$ from (4), (1), (15), and given $\alpha_{k^{\prime},n}$, $\forall n,k^{\prime}\in\mathcal{K}_{n}\setminus\{k\}$; update and save $\boldsymbol{\alpha}^{(i)}$.
@@
8 if $\Big{|}\boldsymbol{\alpha}^{(i)}\!-\!\boldsymbol{\alpha}^{(i-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{1}}$ then set $i\!=\!i\!+\!1$, and goto 3.
9 Compute $w_{1}^{(j)}$ and $w_{2}^{(j)}$; save $\mathbf{W}^{(j)}$.
10 if $\Big{|}\mathbf{W}^{(j)}\!-\!\mathbf{W}^{(j-1)}\Big{|}\!>\!\boldsymbol{\varepsilon_{2}}$, $\mathbf{W}^{(j)}\preceq 0$, $\mathbf{W}^{(j)}\succeq 1$, or $\exists_{k,n}\alpha_{k,n}^{(i)}<0$ then step up $\eta^{(j)}$, $j\!=\!j\!+\!1$, and goto 1.
Return $\boldsymbol{\alpha}^{*}=\boldsymbol{\alpha}^{(i)}$ and $\mathbf{W}^{*}=\mathbf{W}^{(j)}$.
Algorithm 1 Solution for deriving $\boldsymbol{\alpha}^{*}$ and $\mathbf{W}^{*}$ at the $t$-th service interval.
@@ IV-C Analysis for Large $I_{n}$
We consider $\mathcal{X}_{n}$, $\forall n$, has large enough state space, i.e., $I_{n}\gg 1$. In this regime, we achieve $\underset{I_{n}\to\infty}{\lim}P_{{\rm e},n}=\frac{1}{2}$ from (4). Besides, we rewrite (1) and define ${\mathcal{E}}_{n}^{(k)}$ and ^$\widehat{\mathcal{E}}_{n}^{(k)}$ as below
${\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}_{\ell}}\mathcal{E}_{n,\ell}}{\alpha_{k,n}},~{}\,\widehat{\mathcal{E}}_{n}^{(k)}=\dfrac{\sum_{\ell:k\in\mathcal{R}^{\prime}_{\ell}}\mathcal{E}_{n,\ell}}{1-\alpha_{k,n}}.$
Thereby, using (15), $\alpha^{*}_{k,n}$, $\forall k,n$, is derived as follows
$\alpha_{k,n}^{*}=\dfrac{\eta\overline{\mathcal{S}}_{n}{\rm{EUU}}_{\rm min}\Big{(}1\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}-2w_{1}g_{1}\!\!\left(\dfrac{\mathcal{F}_{1}}{2^{(|\widehat{\mathcal{A}}_{t}|+1)}}\right)}{\eta\overline{\mathcal{S}}_{n}{\rm{EUU}}_{\rm min}\Big{(}{\mathcal{E}}_{n}^{(k)}\!-\widehat{\mathcal{E}}_{n}^{(k)}\Big{)}\!}.$ (17)
From (17), we infer that $\alpha_{k,n}^{*}$ increases by (i) decreasing the average usefulness of updates to tackle the high contamination of useful updates over the channels, (ii) enlarging the set of queried attributes, and (iii) increasing the ${\rm{EUU}}_{\rm min}$ constraint.